OpenStudy (anonymous):

solve the following equation for x z=1/28(5y-11x)= plezz help

OpenStudy (anonymous):

z=1/28(y) - 11/28(x) z - (1/28)(y) = (-11/28)(x) [z-(1/28)(y)] / (-11/28) = x x = [ z - (1/28) (y) ] * (-28/11) x = -(28/11)(z) + (1/11)(y)